OpenStudy (anonymous):

What intervals is f'(x) = (2x(x^2-25)-2x(x^2))/((x^2-25)^2) increasing and decreasing?

OpenStudy (zehanz):

If you simplify f', you get:\[f'(x)=\frac{ 2x(x^2-25)-2x*x^2 }{ (x^2-25)^2 }=\frac{ 2x(x^2-25-x^2) }{ (x^2-25)^2 }\]\[f'(x)=-\frac{ 50x }{ (x^2-25)^2 }\]We have to differentiate f'again and solve f''(x) = 0 to see if where f' is increasing and decreasing:\[f''(x)=-\frac{ 50(x^2-25)^2 -50x*2(x^2-25)*2x }{ (x^2-25)^4 }\]Simplify:\[f''(x)=-\frac{ 50(x^2-25)(x^2-25-4x^2) }{ (x^2-25)^4 }\]\[f''(x)=-\frac{ 50(-3x^2-25) }{ (x^2-25)^3 }\]If you try to solve f''(x) = 0 you can see that 1. -3x²-25 = 0 2. x = 5 or x = -5 also have influence on the sign of f'', so make a sign scheme with the zeroes of th nominator AND the denominator.

OpenStudy (zehanz):

In fact, I hope I made a mistake somewhere, because I don't like the nominator ;) Try to differentiate yourself (you need the Quotient Rule and the Chain Rule)